OpenStudy (anonymous):

(2p with a negative 1 exponent) cubed

jimthompson5910 (jim_thompson5910):

\[\large (2p^{-1})^3\] \[\large 2^3(p^{-1})^3\] \[\large 2^3p^{-1*3}\] \[\large 2^3p^{-3}\] \[\large \frac{8}{p^{3}}\] So \[\large (2p^{-1})^3=\frac{8}{p^{3}}\]